# Unveiling Transformer Perception by Exploring Input Manifolds

### v1

@@ 5 Related work
Our work relates to embedding space exploration literature, and has at least one collateral applications in the XAI domain, namely producing feature importance-based explanations.
Embedding space exploration.
Works dealing with embedding space exploration mostly focus on the study of specific properties of the embedding space of Transformers, especially in NLP. For instance,  Cai et al. [5] challenge the idea that the embedding space is inherently anisotropic [10] discovering local isotropy, and find low-dimensional manifold structures in the embedding space of GPT and BERT.  Biś et al. [3] argue that the anisotropy of the embedding space derives from embeddings shifting in common directions during training. In the field of CV,  Vilas et al. [21] map internal representations of a ViT onto the output class manifold, enabling the early identification of class-related patches and the computation of saliency maps on the input image for each layer and head. Applying Singular Value Decomposition to the Jacobian matrix of a ViT,  Salman et al. [17] treat the input space as the union of two subspaces: one in which image embedding doesn’t change, and another one for which it changes. Except for the last one, all the aforementioned approaches rely on data samples. By studying the inverse image of the model, instead, we can do away with data samples.
Feature importance-based explanations.
Feature importance is a measure of the contribution of each data feature to a model prediction. In the context of Computer Vision and Natural Language Processing, it amounts to giving a weight to pixels (or patches of pixels) in an image and tokens in a piece of text, respectively. In recent years, much research has focused on Transformers in both CV and NLP. Most approaches are based on the attention mechanism of the Transformer architecture. Abnar and Zuidema [1] quantify the overall attention of the output on the input by computing a linear combination of layer attentions (Attention Rollout) or applying a maximum flow algorithm (Attention Flow). To overcome the limitations [4] of attention-based methods, Hao et al. [11] use the concept of attribution, which is obtained by multiplying attention matrices by the integrated gradient of the model with respect to them. Chefer et al. [6] propose the Relevancy metric to generalize attribution to bi-modal and encoder-decoder architectures. Other methods are perturbation-based, where perturbations of input data are used to record any change in the output and draw a saliency map on the input. In order to overcome the main issue with such methods, i.e. the generation of outlier inputs, Englebert et al. [9] apply perturbations after the position encoding of the patches. In contrast with these methods, ours does not need arbitrary perturbations of inputs, and considers all parameters of the model, not only the attention query and key matrices.
@@ 6 Conclusions
Our exploration of the Transformer architecture through a theoretical framework grounded in Riemannian Geometry led to the application of our two algorithms, SiMEC and SiMExp, for examining equivalence classes in the Transformers’ input space. We demonstrated how the results of these exploration methods can be interpreted in a human-readable form and conducted preliminary investigations into their potential applications. Notably, our methods show promise for ranking feature importance and generating alternative prompts within the same or different equivalence classes.
Future research directions include expanding our experimental results and delving deeper into the potential of our framework for controlled input generation within an equivalence class. This application holds significant promise for enhancing the explainability of Transformer models’ decisions and for addressing issues related to bias and hallucinations.

### v2

@@ 5 Related work
Works dealing with embedding space exploration mostly focus on the study of specific properties of the embedding space of Transformers, especially in NLP. For instance,  Cai et al. [6] challenge the idea that the embedding space is inherently anisotropic [9] discovering local isotropy, and find low-dimensional manifold structures in the embedding space of GPT and BERT.  Biś et al. [4] argue that the anisotropy of the embedding space derives from embeddings shifting in common directions during training. In the field of Computer Vision,  Vilas et al. [23] map internal representations of a ViT onto the output class manifold, enabling the early identification of class-related patches and the computation of saliency maps on the input image for each layer and head. Applying Singular Value Decomposition to the Jacobian matrix of a ViT,  Salman et al. [17] treat the input space as the union of two subspaces: one in which image embedding doesn’t change, and another one for which it changes. Except for the last one, all the aforementioned approaches rely on data samples. By studying the inverse image of the model, instead, we can do away with data samples. The idea of applying Riemannian geometry to capture geometric information about the input manifold of a neural network building a foliation in equivalence classes has also been explored in [10, 21, 22] in the case of simple architectures. In these works a foliation of the data domain is obtained by means of the pullback of a variation of the Fisher information matrix for classifier networks with ReLU and softmax activation functions, with applications to knowledge transfer and the study of adversarial attacks.
In contrast to these works, we apply Riemannian geometry techniques to study the embedding space of transformers, computing the pullback of the metric of the output space, and we address the further problem of interpreting the output of the exploration process. Furthermore, our algorithms explore the embedding space dynamically, with a non-fixed choice of the integration step $\delta$.
@@ 6 Conclusions
Our exploration of the Transformer architecture through a theoretical framework grounded in Riemannian geometry led to the application of our two algorithms, SiMEC and SiMExp, for examining equivalence classes in the Transformers’ input space. In particular, our method enables two complementary exploration strategies, one for retrieving input instances that produce the same class probability distribution as the original instance, the other for discovering instances that yield a different class probability distribution. We demonstrated how the results of these exploration methods can be interpreted in a human-readable form and how the exploration outputs can be used to generate alternative input data.
Future research directions include delving deeper into the potential of our framework for controlled input generation within an equivalence class. Our goal is to investigate how, in the XAI scenario, our framework can facilitate local and task-agnostic explainability methods applicable to Computer Vision (CV) and Natural Language Processing (NLP) tasks, among others. In particular, we see our methods as potential approaches to investigate Transformers’ sensitivity and explainability with respect to input data features. In future applications to large-scale architectures where the dimension of the embedding space is of order $10^{3}-10^{4}$ [5, 19], we also plan to improve the scalability of the SiMEC and SiMExp algorithms, e.g. making use of partial decompositions.
@@
This work is partially supported by PNRR-NGEU program under MUR 118/2023.